# Solving Optimal Power Flow using a Variational Quantum Approach

### v1

@@ 8 Numerical tests
9 Conclusions
This work has proposed a doubly variational quantum approach to solve the OPF, which is a large-scale nonconvex QCQP. Primal/dual OPF variables are modeled by scaling the state and the PMF corresponding to two PQCs. These two PQCs are trained using exclusively the gradient estimates of the Lagrangian function. The training procedure aims to seek an approximate stationary point of the Lagrangian function. By adopting the XBM method and leveraging the sparsity patterns of power systems, the OPF observables are judiciously permuted, resulting in a banded form so that the Lagrangian function can be measured efficiently. Numerical tests on the IEEE 57-node benchmark power system using PennyLane’s quantum simulator have demonstrated that: i) the EG iteration performs markedly better than the PD iteration; ii) solving (QCQPθ) on PQCs offers smaller errors in terms of the generator setpoint and dual variables, as well as smaller constraint violations compared to solving (QCQP) on a classical computer using PD; and iii) the Q-OPF found generator setpoints and dual variables with small errors, and under the generator setpoints obtained by the Q-OPF, the network constraints are nearly satisfied.
@@ 9 Conclusions
10 Acknowledgments
VK and TL acknowledge support from the US National Science Foundation under grant no. 2412947, and the US Office of Naval Research under grant N000142412614. MMW acknowledges support from the National Science Foundation under grant no. 2329662 and from the Cornell School of Electrical and Computer Engineering.
range
pages10 rangepages16 rangepages1 rangepages10 rangepages16 rangepages6 rangepages3 rangepages1 rangepages1 rangepages15 rangepages19 rangepages1 rangepages1 rangepages1 rangepages1 rangepages1 rangepages1 rangepages1 rangepages1 rangepages11 rangepages10 rangepages24 rangepages6 rangepages1 rangepages8 rangepages30 rangepages24 rangepages1 rangepages1 rangepages1 rangepages12 rangepages36 rangepages11 rangepages10 rangepages9 rangepages1 rangepages16 rangepages1 rangepages5 rangepages7 rangepages4 rangepages1 rangepages1 rangepages8 rangepages1 rangepages1
@@ A.1 Power system modeling
The power injection $(p_{n},q_{n})$ consists of a dispatchable constituent $(p^{g}_{n},q^{g}_{n})$ and an inflexible constituent $(p^{d}_{n},q^{d}_{n})$ such that $p_{n}=p^{g}_{n}-p^{d}_{n}$ and $q_{n}=q^{g}_{n}-q^{d}_{n}$. The former represents the power dispatch of a generator or a flexible load placed at node $n$, and the latter expresses the inelastic load supplied at node $n$. Let nodes with dispatchable power injections belong to a subset $\mathcal{N}_{g}\subseteq\mathcal{N}$. The remaining nodes hosted inelastic loads form the subset $\mathcal{N}_{\ell}=\mathcal{N}\backslash\mathcal{N}_{g}$. Node $n=1$ is referred to as the angle reference node, under which it follows $\mathbf{v}^{\dagger}\mathbf{e}_{1}\mathbf{e}_{1}^{\top}\mathbf{v}=1$, where $\mathbf{e}_{1}$ is the first column of the identity. To simplify the exposition, each node in $\mathcal{N}_{g}$ is assumed to host only one dispatchable unit. Given the inflexible loads $\{p^{d}_{n},q^{d}_{n}\}_{n\in\mathcal{N}}$, the OPF’s objective is to minimize the dispatch cost of generators and flexible load while respecting the resource and network limits. The OPF is formulated as the ensuing quadratically constrained quadratic program (QCQP) over nodal voltages and power of generators:
$\min\leavevmode\nobreak\$$\leavevmode\nobreak\ \sum_{n\in\mathcal{N}_{g}}c_{n}p^{g}_{n}$ (44a)
$\mathrm{over}\leavevmode\nobreak\$$\leavevmode\nobreak\ \mathbf{v}\in\mathbb{C}^{N},\{p^{g}_{n},q^{g}_{n}\}_{n\in\mathcal{N}_{g}}$
$\mathrm{s.to}\leavevmode\nobreak\$$\leavevmode\nobreak\ \mathbf{v}^{\dagger}\mathbf{M}_{p_{n}}\mathbf{v}=p^{g}_{n}-p^{d}_{n}$ $\forall n\in\mathcal{N}_{g}$ (44b)
$\leavevmode\nobreak\ \mathbf{v}^{\dagger}\mathbf{M}_{q_{n}}\mathbf{v}=q^{g}_{n}-q^{d}_{n}$ $\forall n\in\mathcal{N}_{g}$ (44c)
$\leavevmode\nobreak\ \mathbf{v}^{\dagger}\mathbf{M}_{p_{n}}\mathbf{v}=-p^{d}_{n}$ $\forall n\in\mathcal{N}_{\ell}$ (44d)
$\leavevmode\nobreak\ \mathbf{v}^{\dagger}\mathbf{M}_{q_{n}}\mathbf{v}=-q^{d}_{n}$ $\forall n\in\mathcal{N}_{\ell}$ (44e)
$\leavevmode\nobreak\ \underaccent{\bar}{p}^{g}_{n}\leq\mathbf{v}^{\dagger}\mathbf{M}_{p_{n}}\mathbf{v}+p^{d}_{n}\leq\bar{p}^{g}_{n}$ $\forall n\in\mathcal{N}_{g}$ (44f)
$\leavevmode\nobreak\ \underaccent{\bar}{q}^{g}_{n}\leq\mathbf{v}^{\dagger}\mathbf{M}_{q_{n}}\mathbf{v}+q^{d}_{n}\leq\bar{q}^{g}_{n}$ $\forall n\in\mathcal{N}_{g}$ (44g)
$\leavevmode\nobreak\ \underaccent{\bar}{v}_{n}\leq\mathbf{v}^{\dagger}\mathbf{M}_{v_{n}}\mathbf{v}\leq\bar{v}_{n}$ $\forall n\in\mathcal{N}$ (44h)
$\leavevmode\nobreak\ \mathbf{v}^{\dagger}\mathbf{e}_{1}\mathbf{e}_{1}^{\top}\mathbf{v}=1$ (44i)
$\leavevmode\nobreak\ \mathbf{v}^{\dagger}\mathbf{M}_{i_{nm}}\mathbf{v}\leq\bar{i}_{nm}$ $\forall(n,m)\in\mathcal{E}$ (44j)
@@ A.3 Proof of Lemma 3
Quantum observables are trigonometric functions of $(\boldsymbol{\theta},\boldsymbol{\phi})$. Moreover, the Lagrangian function $\mathcal{L}(\mathbf{z})$ is quadratic in $(\alpha,\beta)$. Therefore, the operator $\mathbf{g}(\mathbf{z})$ is continuously differentiable with respect to $\mathbf{z}$. Let $\mathbf{J}(\mathbf{z})\coloneqq\nabla_{\mathbf{z}}\mathbf{g}(\mathbf{z})$ denote the Jacobian matrix of $\mathbf{g}(\mathbf{z})$. If the spectral norm of $\mathbf{J}(\mathbf{z})$ is upper bounded by a constant $L$, then $\mathbf{g}(\mathbf{z})$ is $L$-Lipschitz continuous [61, Lemma 2.6]. Based on matrix norm inequalities, the spectral norm of $\mathbf{J}(\mathbf{z})$ can be bounded as:
@@
According to (46), the Lipschitz constant $L$ of $\mathbf{g}(\mathbf{z})$ is the maximum of the right-hand sides of (A.3)–(A.3). If $\alpha\geq 4$, the right-hand side (RHS) of (A.3) is larger than the RHS of (A.3). The condition $\alpha\geq 4$ holds trivially if the number of power system nodes is $N>32$. Moreover, if $\beta\geq 2$, the RHS of (A.3) is larger than that of (A.3). The condition $\beta\geq 2$ is expected to occur for problems with many constraints. Therefore, we are left with the RHS of (A.3) and (A.3), whose maximum can be compactly expressed as
@@ A.4 Proof of Lemma 4
We first upper bound the variance of the sample-average estimators of the three terms in (9). Note that the three terms in (9) can be decomposed into quantum expectations across $c$. This is obvious for $F_{0}(\boldsymbol{\theta})$ and $G(\boldsymbol{\phi})$. Regarding $F(\boldsymbol{\theta},\boldsymbol{\phi})$, we obtain from (6.3):

### v2

@@ VIII Numerical Tests
IX Conclusions
This work proposes a doubly variational quantum approach to solve QCQP/SDPs, where the observables are induced by underlying sparse graphs. Primal/dual variables are modeled by scaling the state and the PMF corresponding to two PQCs. These two PQCs are trained using exclusively the gradient estimates of the Lagrangian function. The training procedure aims to seek an approximate stationary point of the Lagrangian function. By adopting the XBM method and leveraging graph sparsity, the related observables are judiciously permuted into a banded form, enabling efficient evaluation of the Lagrangian function. The proposed algorithmic framework is applied to the OPF problem. Numerical tests on the IEEE 57-node benchmark power system using PennyLane’s quantum simulator have demonstrated that: i) the EG iteration performs markedly better than the PD iteration; ii) solving (QCQPθ) on PQCs offers smaller errors in terms of the generator setpoint and dual variables, as well as smaller constraint violations compared to solving (QCQP) on a classical computer using PD; and iii) the Q-OPF found generator setpoints and dual variables with small errors, and under the generator setpoints obtained by the Q-OPF, the network constraints are nearly satisfied.
@@ IX Conclusions
Acknowledgements.
VK and TVL acknowledge support from the US National Science Foundation under grant no. 2412947, and the US Office of Naval Research under grant N000142412614. MMW acknowledges support from the National Science Foundation under grant no. 2329662 and from the Cornell School of Electrical and Computer Engineering.
@@ A.1 Power system modeling
The power injection $(p_{n},q_{n})$ consists of a dispatchable constituent $(p^{g}_{n},q^{g}_{n})$ and an inflexible constituent $(p^{d}_{n},q^{d}_{n})$ such that $p_{n}=p^{g}_{n}-p^{d}_{n}$ and $q_{n}=q^{g}_{n}-q^{d}_{n}$. The former represents the power dispatch of a generator or a flexible load placed at node $n$, and the latter expresses the inelastic load supplied at node $n$. Let nodes with dispatchable power injections belong to a subset $\mathcal{N}_{g}\subseteq\mathcal{N}$. The remaining nodes hosted inelastic loads form the subset $\mathcal{N}_{\ell}=\mathcal{N}\backslash\mathcal{N}_{g}$. Node $n=1$ is referred to as the angle reference node, under which it follows $\mathbf{v}^{\dagger}\mathbf{e}_{1}\mathbf{e}_{1}^{\top}\mathbf{v}=1$, where $\mathbf{e}_{1}$ is the first column of the identity. To simplify the exposition, each node in $\mathcal{N}_{g}$ is assumed to host only one dispatchable unit. Given the inflexible loads $\{p^{d}_{n},q^{d}_{n}\}_{n\in\mathcal{N}}$, the OPF’s objective is to minimize the dispatch cost of generators and flexible load while respecting the resource and network limits. The OPF is formulated as the ensuing QCQP over nodal voltages and power of generators:
$\min\penalty 10000\$$\penalty 10000\ \sum_{n\in\mathcal{N}_{g}}c_{n}p^{g}_{n}$ (45a)
$\mathrm{over}\penalty 10000\$$\penalty 10000\ \mathbf{v}\in\mathbb{C}^{N},\{p^{g}_{n},q^{g}_{n}\}_{n\in\mathcal{N}_{g}}$
$\mathrm{s.to}\penalty 10000\$$\penalty 10000\ \mathbf{v}^{\dagger}\mathbf{M}_{p_{n}}\mathbf{v}=p^{g}_{n}-p^{d}_{n}$ $\forall n\in\mathcal{N}_{g}$ (45b)
$\penalty 10000\ \mathbf{v}^{\dagger}\mathbf{M}_{q_{n}}\mathbf{v}=q^{g}_{n}-q^{d}_{n}$ $\forall n\in\mathcal{N}_{g}$ (45c)
$\penalty 10000\ \mathbf{v}^{\dagger}\mathbf{M}_{p_{n}}\mathbf{v}=-p^{d}_{n}$ $\forall n\in\mathcal{N}_{\ell}$ (45d)
$\penalty 10000\ \mathbf{v}^{\dagger}\mathbf{M}_{q_{n}}\mathbf{v}=-q^{d}_{n}$ $\forall n\in\mathcal{N}_{\ell}$ (45e)
$\penalty 10000\ \underaccent{\bar}{p}^{g}_{n}\leq\mathbf{v}^{\dagger}\mathbf{M}_{p_{n}}\mathbf{v}+p^{d}_{n}\leq\bar{p}^{g}_{n}$ $\forall n\in\mathcal{N}_{g}$ (45f)
$\penalty 10000\ \underaccent{\bar}{q}^{g}_{n}\leq\mathbf{v}^{\dagger}\mathbf{M}_{q_{n}}\mathbf{v}+q^{d}_{n}\leq\bar{q}^{g}_{n}$ $\forall n\in\mathcal{N}_{g}$ (45g)
$\penalty 10000\ \underaccent{\bar}{v}_{n}\leq\mathbf{v}^{\dagger}\mathbf{M}_{v_{n}}\mathbf{v}\leq\bar{v}_{n}$ $\forall n\in\mathcal{N}$ (45h)
$\penalty 10000\ \mathbf{v}^{\dagger}\mathbf{e}_{1}\mathbf{e}_{1}^{\top}\mathbf{v}=1$ (45i)
$\penalty 10000\ \mathbf{v}^{\dagger}\mathbf{M}_{i_{nm}}\mathbf{v}\leq\bar{i}_{nm}$ $\forall(n,m)\in\mathcal{E}$ (45j)
@@ A.3 Proof of Lemma 3
Quantum observables are trigonometric functions of $(\boldsymbol{\theta},\boldsymbol{\phi})$. Moreover, the Lagrangian function $\mathcal{L}(\mathbf{z})$ is quadratic in $(\alpha,\beta)$. Therefore, the operator $\mathbf{g}(\mathbf{z})$ is continuously differentiable with respect to $\mathbf{z}$. Let $\mathbf{J}(\mathbf{z})\coloneqq\nabla_{\mathbf{z}}\mathbf{g}(\mathbf{z})$ denote the Jacobian matrix of $\mathbf{g}(\mathbf{z})$. If the spectral norm of $\mathbf{J}(\mathbf{z})$ is upper bounded by a constant $L$, then $\mathbf{g}(\mathbf{z})$ is $L$-Lipschitz continuous [22, Lemma 2.6]. Based on matrix norm inequalities, the spectral norm of $\mathbf{J}(\mathbf{z})$ can be bounded as:
@@
According to (47), the Lipschitz constant $L$ of $\mathbf{g}(\mathbf{z})$ is the maximum of the right-hand sides of (A.3)–(A.3). If $\alpha\geq 4$, the right-hand side (RHS) of (A.3) is larger than the RHS of (A.3). The condition $\alpha\geq 4$ holds trivially if the problem size is $N>32$. Moreover, if $\beta\geq 2$, the RHS of (A.3) is larger than that of (A.3). The condition $\beta\geq 2$ is expected to occur for problems with many constraints. Therefore, we are left with the RHS of (A.3) and (A.3), whose maximum can be compactly expressed as
@@ A.4 Proof of Lemma 4
We first upper bound the variance of the sample-average estimators of the three terms in (10). Note that the three terms in (10) can be decomposed into quantum expectations across $c$. This is obvious for $F_{0}(\boldsymbol{\theta})$ and $G(\boldsymbol{\phi})$. Regarding $F(\boldsymbol{\theta},\boldsymbol{\phi})$, we obtain from (V.3):